OpenStudy (anonymous):

Evaluate the indefinite integral:

OpenStudy (anonymous):

\[\int\limits_{?}^{?}x^4\sqrt{12+ x^5}sx\]

OpenStudy (anonymous):

ignore the question marks

OpenStudy (anonymous):

Okay , you can take the following:\[u= x^5 + 12\]\[du = 5x^4dx\] now you'll get the following: \[=1/5 \int\limits_{}^{}\sqrt{u} du\] You can proceed from here ^_^

OpenStudy (anonymous):

if you still can't solve it, let me know :)

OpenStudy (anonymous):

yeah i'm still kind of stuck :/

OpenStudy (anonymous):

\[1/5\int\limits_{}^{}\sqrt{u} du =1/5\int\limits_{}^{}^{}^{} u ^{1/2} = 1/5*2/3u ^{3/2} + C \] Now plug in u.

OpenStudy (anonymous):

lol, thank you DE ^_^